# Learning-based Multi-View Stereo: A Survey

### v1

@@ I-E Datasets and Benchmarks
Tab. I is a brief summary of commonly used public MVS datasets and benchmarks for training and evaluation.
TABLE I: An overview of commonly used datasets and benchmarks for learning-based Multi-View Stereo.
Dataset Provided Ground Truth1 Synthetic Online Evaluation
Camera pose Depth Map Point Cloud Mesh Benchmark Target
ScanNet [26] Depth Map / Mesh
7-Scenes [27] Depth Map
DTU [21] Point Cloud
Tanks and Temples [22] Point Cloud
ETH3D [23] Point Cloud
BlendedMVS [24] Depth Map
  • 1
    For datasets with online benchmark, the point cloud ground truth of test set is not released.
ScanNet: ScanNet [26] is a large RGB-D dataset that contains 1613 indoor scenes with ground-truth camera poses, depth maps, surface reconstruction and semantic segmentation labels. Online MVS methods [25, 31, 35] mainly use ScanNet for training and testing.
7-Scenes: 7-Scenes dataset [27] is a RGB-D dataset captured in indoor scenes with a handheld Kinect RGB-D camera. Since it is relatively small, 7-Scenes is usually used to test the generalization performance of the models trained on ScanNet [26] without finetuning.
DTU: DTU dataset [21] is an object-centric MVS dataset collected under well-controlled laboratory conditions with known accurate camera trajectory. It contains 128 scans with 49 or 64 views under 7 different lighting conditions. Since DTU dataset officially provides scanned ground truth point clouds instead of depth maps, it is required to generate mesh models with surface reconstruction, e.g., screened Poisson surface reconstruction algorithm [67], and then render depth maps [15] for training.
Tanks and Temples: Tanks and Temples [22] is a large-scale benchmark captured in more complex real indoor and outdoor scenarios. It is divided into intermediate and advanced sets. Different scenes have different scales, surface reflection and exposure conditions. It is usually used to test the generalization performance. Note that Tanks and Temples does not provide ground truth camera parameters, which are usually estimated with Structure-from-Motion methods such as COLMAP [17] or OpenMVG [56].
ETH3D: ETH3D benchmark [23] contains 25 large-scale scenes, including indoor and outdoor, with high-resolution RGB images. The scenes typically contain many low-textured regions and non-Lambertian surfaces, e.g., white walls and reflective floor. In addition, the images are usually sparse and have strong viewpoint variations and occlusions. Therefore, ETH3D is considered as a very challenging benchmark. Similar to Tanks and Temples, ETH3D is usually used to test the generalization performance.
[…]
I-F Evaluation Metrics
@@
As mentioned in Tab. I, based on the ground truth, e.g., depth maps or point clouds, evaluation metrics can be categorized into 2D metrics and 3D metrics.
I-F1 2D Metrics
@@
2D depth-based metrics are commonly used by online MVS methods [25, 31] to evaluate the depth maps [68, 69]. We list the commonly used metrics as follows: mean absolute depth error (Abs), mean absolute relative depth error (Abs Rel), mean absolute inverse depth error (Abs Inv), and inlier ratio with threshold 1.25 ($\delta<1.25$):
[…]
where $d(\mathbf{p})$ denotes the ground truth depth, $\hat{d}(\mathbf{p})$ denotes the estimation, $N$ denotes the number of pixels with valid depth measurements and $\mathbbm{1}[\cdot]$ denotes the indicator function.
I-F2 3D Metrics
@@
3D point cloud evaluation is widely used by offline MVS methods [15, 37]. Note that the reconstructed point clouds should be aligned to ground truth point clouds before evaluation, e.g., by Iterative Closest Point (ICP), if their extrinsics are differently calibrated.
Precision/Accuracy: Precision/Accuracy measures the percentage of predicted points that can be matched to the ground truth point cloud. Considering a point $\mathbf{P}_{p}$ in the predicted point cloud, it is considered to have a good match in the ground truth point cloud $\{\mathbf{P}_{g}\}$ if
$\|\mathbf{P}_{p}-\mathop{\arg\min}\limits_{\mathbf{P}\in\{\mathbf{P}_{g}\}}\|\mathbf{P}-\mathbf{P}_{p}\|_{2}\|_{2}\leq\lambda,$ (11)
where $\lambda$ is a scene-dependent threshold assigned by datasets. Usually $\lambda$ is set to a large value for large-scale scenes. Note that in some datasets, instead of being measured by percentage [22, 23], precision/accuracy is measured by mean or median absolute distance [21].
Recall/Completeness: Recall/Completeness measures the percentage of ground truth points that can be matched to the predicted point cloud. For a point $\mathbf{P}_{g}$ in the ground truth point cloud, it is considered a good match in the predicted point cloud $\{\mathbf{P}_{p}\}$ if
$\|\mathbf{P}_{g}-\mathop{\arg\min}\limits_{\mathbf{P}\in\{\mathbf{P}_{p}\}}\|\mathbf{P}-\mathbf{P}_{g}\|_{2}\|_{2}\leq\lambda.$ (12)
Similar to precision/accuracy, recall/completeness is sometimes measured by the percentage of points [22, 23] and sometimes measured by mean or median absolute distance [21], whose definition is similar to Chamfer distance.
F-Score: The two aforementioned metrics measure the accuracy and completeness of predicted point clouds. However, each of these metrics alone cannot present the overall performance since different MVS methods adopt different assumptions. A stronger assumption usually leads to higher accuracy but lower completeness. If only precision/accuracy is reported, it would favor MVS algorithms that only include estimated points of high certainty. On the other hand, if only recall/completeness is reported it would favor MVS algorithms that include everything, regardless of accuracy. Therefore, F-score, an integrated metric, is introduced [22, 23]. F-score is the harmonic mean of precision and recall. It is sensitive to extremely small values and tends to get more affected by smaller values, which means that F-score does not encourage imbalanced results. However, in most cases, F-score still suffers from unfairness due to the limitations of ground truth, e.g., sparse and incomplete point clouds may penalize for filling in the areas that are not present in the ground truth [47].
II Supervised Method with Depth Estimation
@@
This section mainly introduces supervised learning-based MVS methods with depth estimation. A typical depth map-based MVS pipeline mainly consists of the feature extraction (Sec. II-A), cost volume construction (Sec. II-B), cost volume regularization (Sec. II-C) and depth estimation (Sec. II-E). The pipelines of MVDepthNet [25] and MVSNet [15] are illustrated in Fig. 4 and Fig. 5, which are the representative methods of online and offline MVS methods respectively.
II-A Feature Extraction
@@
Considering efficiency, most methods use simple CNN structures to extract deep features from images, e.g., ResNet [70], U-Net [71] and FPN [72]. For online MVS methods, feature extraction networks are usually used in line with the real-time operation goal. DeepVideoMVS [31] combines MNasNet [73], which is lightweight and has low latency, with FPN. SimpleRecon [35] proposes utilizing the first two blocks from ResNet18 [70] and an EfficientNet-v2 [74] encoder, which maintains efficiency and yields a sizeable improvement in depth map accuracy. For offline MVS methods, MVSNet [15] uses a stacked eight-layer 2D CNN to extract deep features for all the images. Coarse-to-fine methods further extract multi-scale features for estimation on multiple scales with FPN [32, 39, 40] or multi-scale RGB images [75, 76]. Recently, many following works have been paying more attention to feature extraction to improve the representation power of deep features. [77, 78, 79] introduce deformable convolutions to adaptively learn receptive fields for areas with varying richness of texture. CDS-MVSNet [80] uses a dynamic scale network, CDSFNet, to extract the discriminative features by analyzing the normal curvature of the image surface. GeoMVSNet [76] proposes a geometry fusion network that fuses the image features with coarse depth maps. With FPN as the main feature extractor, WT-MVSNet [81] and MVSFormer [82], further introduces Vision Transformers [83, 84, 85] for image feature enhancement. ET-MVSNet [86] integrates the Epipolar Transformer into the feature extraction, which performs non-local feature aggregation on the epipolar lines.
II-B Cost Volume Construction
@@
For both online and offline MVS methods, the cost volume is constructed with the plane sweep algorithm as discussed in Sec. I-C.
II-B1 Online MVS
@@
To reduce computation and improve the efficiency for online applications, online MVS methods usually construct 3D cost volumes $\mathbf{C}\in\mathbb{R}^{H\times W\times D}$, which store a single value as matching cost for each pixel $\mathbf{p}$ and depth sample $d_{i}$. MVDepthNet [25] and GP-MVS [58] compute the per-pixel intensity difference between the reference and each source view as matching cost. If there is more than one source view, the cost volumes are averaged. Instead of intensity difference, Neural RGB-D [87] computes deep feature difference. DeepVideoMVS [31] and MaGNet [88] calculate the cost from pixel-wise correlation between the reference feature and the warped source feature. In addition to the dot product between reference image features and warped source image features, SimpleRecon [35] further adds metadata, e.g., ray direction and relative pose, into the 4D cost volume. Then an MLP is used to reduce the dimension of 4D cost volume to a 3D cost volume.
II-B2 Offline MVS
@@
Offline MVS methods mainly focus on reconstructing high-quality dense geometry with high-resolution images. To encode more matching information and improve the quality, offline methods [15, 32] usually construct 4D cost volumes $\mathbf{C}\in\mathbb{R}^{H\times W\times D\times C}$, where each pixel $\mathbf{p}$ and depth sample $d_{i}$ is associated with a matching cost of dimension $C$. Since there may exist an arbitrary number of source views and serious occlusion [23], robustly aggregating matching information from all the source views is an important step. MVSNet [15] proposes a variance-based cost metric as follows:
@@
II-C Cost Volume Regularization
@@
II-C1 Online MVS
A 2D encoder-decoder architecture is commonly used to aggregate information. MVDepthNet [25] and GP-MVS [58] concatenate the reference image and cost volume and send them to an encoder-decoder architecture with skip connections. In addition, GP-MVS [58] uses Gaussian process to fuse information from previous views. Neural-RGBD [87] accumulates depth probability volumes over time with a Bayesian filtering framework to effectively reduce depth uncertainty and improve robustness and temporal stability. DeepVideoMVS [31] applies 2D U-Net [71] on the cost volume and adds skip connections between the image encoder and cost volume encoder at all resolutions. It further uses ConvLSTM [93] to propagate past information with a small overhead of computation time and memory consumption. Long et al. [94] use MatchNet to regularize the cost volume and ContextNet to learn 2D context information from the reference view. By concatenating the cost volume and context feature, a hybrid volume is fed into the Epipolar Spatio-Temporal transformer to aggregate temporal information. SimpleRecon [35] fuses multi-scale image features, extracted from the pretrained EfficientNetv2 [74], into the cost volume encoder to improve the performance.
@@ II-C1 Online MVS
II-C2 Offline MVS
Among most learning-based offline MVS methods that use 4D cost volumes, there are three main categories to perform cost volume regularization: direct 3D CNN, coarse-to-fine and RNN. Fig. 6 illustrates these three main schemes.
Figure 6: Illustration of typical cost regularization schemes for offline MVS methods that use 4D cost volumes. (a) direct 3D CNN regularization [15, 95, 96, 36] applies a 3D CNN to aggregate contextual information; (b) coarse-to-fine regularization [32, 39, 75, 63] constructs multi-scale cost volumes based on coarse prediction and uses 3D CNN regularization on each scale; (c) RNN regularization [37, 38, 77] sequentially regularizes 2D slices of the cost volume to reduce memory consumption.
@@ II-D Iterative Update
Diverging from conventional approaches, certain methods [40, 41, 101, 102] adopt iterative updates to gradually refine depth maps. Iterative methods introduce a dynamic approach to depth map estimation, enabling iterative refinement of the reconstruction process. By iteratively updating the depth maps based on successive iterations, these methods can progressively improve the accuracy and consistency of the reconstructed 3D scene. This iterative refinement is particularly advantageous in scenarios where initial estimations may be coarse or inaccurate, as it allows for finer adjustments to be made to the depth maps over multiple iterations. Moreover, the ability to control the number of iterations provides users with the flexibility to prioritize either computational efficiency or reconstruction quality, depending on specific application requirements.
@@
Some methods [40, 103] combine iterative PatchMatch [104] with deep learning. PatchMatch algorithm is widely used in many traditional MVS methods [105, 18, 34, 19]. The PatchMatch algorithm mainly consists of: random initialization, propagation of hypotheses to neighbors, and evaluation for choosing best solutions. After initialization, the approach iterates between propagation and evaluation until convergence. Traditional methods usually design fixed patterns for propagation, e.g., propagation of Gipuma [18] as shown in Fig. 7. Recently, PatchmatchNet [40] proposes learned adaptive propagation and cost aggregation modules, which enables PatchMatch to converge faster and deliver more accurate depth maps. Instead of propagating depth samples naively from a static set of neighbors as done traditionally [18], PatchmatchNet adaptively gathers samples from pixels of the same surface with the guidance of image features, shown in Fig. 8. It iteratively updates the depth maps in a coarse-to-fine structure. Without 3D CNN, PatchmatchNet demonstrates a remarkably low computation time and memory than other coarse-to-fine methods [32, 39, 75]. Another PatchMatch-based method, PatchMatch-RL [103], jointly estimates depth, normal and visibility with a coarse-to-fine structure. Considering argmax based hard decisions/sampling of PatchMatch is non-differentiable, PatMatch-RL adopts reinforcement learning in the training process.
Figure 7: Traditional chessboard propagation scheme of Galliani et al. [18]. Planes from a local neighborhood (red points) serve as candidates to update a given pixel (black). (a) Standard pattern with 20 neighbors. (b) Simplified pattern with 8 neighbors.
Figure 8: Adaptive propagation scheme of Wang et al. [40]. Pixels located at the object boundary (yellow) and a textureless region (red) receive depth hypotheses from sampled neighbors (green and orange). (a) Reference image. (b) Fixed sampling locations of classic propagation. (c) Adaptive sampling locations with adaptive propagation. The grayscale image in (b) and (c) is the ground truth depth map.
@@
Recently, RAFT [106] estimates optical flow by iteratively updating a motion field through a GRU and achieves state-of-the-art performance. The idea is further adopted in stereo [107], scene flow [108] and SfM [109]. In MVS, several methods [41, 102, 101] also adopted this approach to enhance efficiency and flexibility. IterMVS [41] proposes a lightweight GRU-based probability estimator that encodes the per-pixel probability distribution of depth in its hidden state. In each iteration, multi-scale matching information is injected to update the pixel-wise depth distribution. CER-MVS [101] and Effi-MVS [102] both embed the RAFT module in a coarse-to-fine structure. With the injection of matching information, the hidden state is updated and outputs a residual that is added to the previous depth estimation. DELS-MVS [110] proposes Epipolar Residual Network to search for the corresponding point in the source image directly along the corresponding epipolar line and follows an iterative manner to narrow down the search space. Based on IterMVS [41], RIAV-MVS [111] iteratively refines index fields with a 3-level GRUs. MaGNet [88] iteratively updates the Gaussian distribution of depth for each pixel with the matching scores. IGEV-MVS [112] iteratively updates a disparity map regressed from geometry encoding cost volumes and pairs of group-wise correlation volumes.
II-E Depth Estimation
II-E1 Online MVS
@@ II-E2 Offline MVS
where $\mathbf{p}$ is the pixel coordinate, $d_{i}$ is $i$-th depth sample and $P(i,\mathbf{p})$ is the predicted depth probability. For coarse-to-fine methods [32, 39, 75], soft argmax is applied on each stage to regress the depth maps. On the other hand, some methods [36, 40] compute the expectation value of inverse depth samples since this sampling is more suitable for complex and large-scale scenes [36]. The depth prediction is computed as:
$\hat{d}(\mathbf{p})=\left(\sum_{i=1}^{D}\frac{1}{d_{i}}\cdot\mathbf{P}(i,\mathbf{p})\right)^{-1}.$ (17)
@@
Comparatively, soft argmax operation corresponds to measuring the distance of the expectation to the ground truth depth. While the expectation can take any continuous value, the measure cannot handle multiple modes in $\mathbf{P}$ and strictly prefers unimodal distributions [41]. The argmax operation corresponds to measuring the Kullback-Leibler divergence between a one-hot encoding of the ground truth and $\mathbf{P}$, but cannot achieve sub-pixel precision [41]. Recently, Wang et al. [41] propose a simple yet effective hybrid strategy that combines regression and classification. Similar to [113], it is robust to multi-modal distributions but also achieves sub-pixel precision. First, similar to R-MVSNet [37], the index $\mathbf{X}(\mathbf{p})$ with the highest probability for pixel $\mathbf{p}$ from probability $\mathbf{P}$ is found. Then the expectation in the local inverse range is computed as the depth estimate $\hat{d}(\mathbf{p})$:
$\hat{d}(\mathbf{p})\!\!=\!\!\left(\!\!\frac{1}{\sum_{j=\mathbf{X}(\mathbf{p})-r}^{\mathbf{X}(\mathbf{p})+r}\mathbf{P}(j,\mathbf{p})}\!\sum_{j=\mathbf{X}(\mathbf{p})-r}^{\mathbf{X}(\mathbf{p})+r}\!\!\frac{1}{d_{j}}\!\cdot\!\mathbf{P}(j,\mathbf{p})\!\!\right)^{\!\!\!-1}\!\!\!\!\!,$ (18)
where $d_{j}$ is the $j$-th depth sample. Concurrently, Peng et al. [114] propose a different strategy to unify regression and classification. They first use classification to get the optimal hypothesis and then regress the proximity for it.
@@
II-F Depth Refinement
Given that the raw depth estimation from MVS may be noisy, refinement is usually used to improve the accuracy. R-MVSNet [37] enforces multi-view photo-consistency to alleviate the stair effect and achieve sub-pixel precision. Point-MVSNet [95] and VA-Point-MVSNet [116] use PointFlow to refine the point cloud iteratively by estimating the residual between the depth of the current iteration and that of the ground truth. Fast-MVSNet [117] adopts an efficient Gauss-Newton layer to optimize the depth map by minimizing the feature residuals. PatchmatchNet [40] refines the final upsampled depth map with a depth residual network [118] and reference image feature.  [41, 102, 101] use the mask upsampling module from RAFT [106] to upsample and refine the depth map to full resolution by computing the weighted sum of depth values in a window based on the mask. Based on a coarse depth map, RayMVSNet [119, 120] aggregates multi-view image features with an epipolar transformer and use a 1D implicit field to estimate the SDF of the sampled points and the location of the zero-crossing point. GeoMVSNet [76] filters the depth map by geometry enhancement in the frequency domain. EPNet [121] uses a hierarchical edge-preserving residual learning module to refine multi-scale depth estimation with image context features.
II-G Confidence Estimation
As discussed in Sec. I-D, photometric confidence is important to filter out unreliable estimations during depth fusion for offline MVS methods. Following MVSNet [15], most offline MVS methods take the probability of the estimation [37, 38, 77] or the probability sum over several samples near the estimation [15, 32, 40] from the probability volume as confidence. Fig. 9 provides a set of RGB images, estimated depth maps, and the visualized confidence. In stereo matching, some methods learn confidence from disparity [122], RGB image [123, 124] or matching costs [125] and obtain confidence scores in $[0,1]$ interval. Motivated by this, some traditional MVS methods [126, 127] based on classical PatchMatch [105] propose to estimate the confidence with deep learning and use this to refine the results from PatchMatch. Recently, IterMVS [41] estimates confidence from the pixel-wise depth probability distributions, which are encoded by the hidden state of a convolutional GRU. A 2D CNN followed by a sigmoid is applied to the hidden state to predict the confidence. DELS-MVS [110] feeds pixel-wise entropy of the partition probabilities, which is computed for each source image, into a confidence network to learn a confidence map. The confidence map is used to guide the fusion of multiple depth maps.
Figure 9: Visualization of confidence on DTU [21]. The regions with low confidence typically correspond to occlusions and low texture.
II-H Loss Function
II-H1 Online MVS

### v2

@@ 2.5 Datasets and Benchmarks
Along with this survey, we also include a brief summary of commonly used public MVS datasets and benchmarks for training and evaluation. For more details, please refer to supplementary.
2.6 Evaluation Metrics
@@
Based on the ground truth, e.g., depth maps or point clouds, evaluation can be categorized into 2D and 3D metrics.
@@
2D Metrics: 2D depth-based metrics are commonly used by online MVS methods [25, 30] to evaluate the depth maps [68, 69]. The common metrics are: mean absolute depth error (Abs), mean absolute relative depth error (Abs Rel) and inlier ratio with threshold 1.25 ($\delta<1.25$). For more details, please refer to supplementary.
@@
3D Metrics: 3D point cloud evaluation is widely used by offline MVS methods [15, 36]. Note that the reconstructed point clouds should be aligned to ground truth point clouds before evaluation, e.g., by Iterative Closest Point (ICP), if their extrinsics are differently calibrated. The metrics are Precision/Accuracy, Recall/Completeness and F-score. For more details, please refer to supplementary.
3 Supervised Method with Depth Estimation
@@
This section mainly introduces supervised learning-based MVS methods with depth estimation. A typical depth map-based MVS pipeline mainly consists of the feature extraction (Sec. 3.1), cost volume construction (Sec. 3.2), cost volume regularization (Sec. 3.3) and depth estimation (Sec. 3.5). The pipelines of MVDepthNet [25] and MVSNet [15] are illustrated in Fig. 3 and Fig. 4, which are the representative methods of online and offline MVS methods respectively.
3.1 Feature Extraction
@@
Considering efficiency, most methods use simple CNN structures to extract deep features from images, e.g., ResNet [70], U-Net [71] and FPN [72]. For online MVS methods, feature extraction networks are usually used in line with the real-time operation goal. DeepVideoMVS [30] combines MNasNet [73], which is lightweight and has low latency, with FPN. SimpleRecon [34] proposes utilizing the first two blocks from ResNet18 [70] and an EfficientNet-v2 [74] encoder, which maintains efficiency and yields a sizeable improvement in depth map accuracy. For offline MVS methods, MVSNet [15] uses a stacked 8-layer 2D CNN to extract deep features for all the images. Coarse-to-fine methods further extract multi-scale features for estimation on multiple scales with FPN [31, 38, 39] or multi-scale RGB images [75, 76]. Recently, many following works have been paying more attention to feature extraction to improve the representation power of deep features. [77, 78, 79] introduce deformable convolutions to adaptively learn receptive fields for areas with varying richness of texture. CDS-MVSNet [80] uses a dynamic scale network, CDSFNet, to extract the discriminative features by analyzing the normal curvature of the image surface. GeoMVSNet [76] proposes a geometry fusion network that fuses the image features with coarse depth maps. With FPN as the main feature extractor, WT-MVSNet [81] and MVSFormer [82], further introduces Vision Transformers [83, 84, 85] for image feature enhancement. ET-MVSNet [86] integrates the Epipolar Transformer into the feature extraction, which performs non-local feature aggregation on the epipolar lines.
3.2 Cost Volume Construction
@@
For both online and offline MVS methods, the cost volume is constructed with the plane sweep algorithm as discussed in Sec. 2.3.
3.2.1 Online MVS
@@
To reduce computation and improve the efficiency for online applications, online MVS methods usually construct 3D cost volumes $\mathbf{C}\in\mathbb{R}^{H\times W\times D}$, which store a single value as matching cost for each pixel $\mathbf{p}$ and depth sample $d_{i}$. MVDepthNet [25] and GP-MVS [60] compute the per-pixel intensity difference between the reference and each source view as matching cost. If there is more than one source view, the cost volumes are averaged. Instead of intensity difference, Neural RGB-D [87] computes deep feature difference. DeepVideoMVS [30] and MaGNet [88] calculate the cost from pixel-wise correlation between the reference feature and the warped source feature. In addition to the dot product between reference image features and warped source image features, SimpleRecon [34] further adds metadata, e.g., ray direction and relative pose, into the 4D cost volume. Then an MLP is used to reduce the dimension of 4D cost volume to a 3D cost volume. DoubleTake [89] further includes depth hint, the depth map rendered from the previously predicted geometry, into the cost volume.
3.2.2 Offline MVS
@@
Offline MVS methods focus on reconstructing high-quality dense geometry with high-resolution images. To encode more matching information and improve the quality, offline methods [15, 31] usually construct 4D cost volumes $\mathbf{C}\in\mathbb{R}^{H\times W\times D\times C}$, where each pixel $\mathbf{p}$ and depth sample $d_{i}$ is associated with a matching cost of dimension $C$. Since there may exist an arbitrary number of source views and serious occlusion [23], robustly aggregating matching information from all the source views is an important step. MVSNet [15] proposes a variance-based cost metric as follows:
@@
3.3 Cost Volume Regularization
@@
3.3.1 Online MVS
A 2D encoder-decoder architecture is commonly used to aggregate information. MVDepthNet [25] and GP-MVS [60] concatenate the reference image and cost volume and send them to an encoder-decoder architecture with skip connections. In addition, GP-MVS [60] uses Gaussian process to fuse information from previous views. Neural-RGBD [87] accumulates depth probability volumes over time with a Bayesian filtering framework to effectively reduce depth uncertainty and improve robustness and temporal stability. DeepVideoMVS [30] applies 2D U-Net [71] on the cost volume and adds skip connections between the image encoder and cost volume encoder at all resolutions. It further uses ConvLSTM [95] to propagate past information with a small overhead of computation time and memory consumption. SimpleRecon [34] fuses multi-scale image features, extracted from the pretrained EfficientNetv2 [74], into the cost volume encoder to improve the performance.
@@ 3.3.1 Online MVS
3.3.2 Offline MVS
Among most learning-based offline MVS methods that use 4D cost volumes, there are three main categories to perform cost volume regularization: direct 3D CNN, coarse-to-fine and RNN.
@@ 3.4 Iterative Update
Diverging from conventional approaches, certain methods [39, 102, 40, 103, 104] adopt iterative updates to gradually refine depth maps. By iteratively updating the depth maps based on successive iterations, these methods can progressively improve the accuracy and consistency of the reconstructed 3D scene. Moreover, the ability to control the number of iterations provides users with the flexibility to prioritize either computational efficiency or reconstruction quality, depending on specific application requirements.
@@
Some methods [39, 102] combine iterative PatchMatch [105] with deep learning. PatchMatch algorithm is widely used in many traditional MVS methods [106, 18, 33, 19]. The PatchMatch algorithm mainly consists of: random initialization, propagation of hypotheses to neighbors, and evaluation for choosing best solutions. After initialization, the approach iterates between propagation and evaluation until convergence. Traditional methods usually design fixed patterns for propagation. Recently, PatchmatchNet [39] proposes learned adaptive propagation and cost aggregation modules, which enables PatchMatch to converge faster and deliver more accurate depth maps. Instead of propagating depth samples naively from a static set of neighbors as done traditionally [18], PatchmatchNet adaptively gathers samples from pixels of the same surface with the guidance of image features. PatchMatch-RL [102], jointly estimates depth, normal and visibility with a coarse-to-fine structure. Considering argmax based hard decisions/sampling of PatchMatch is non-differentiable, PatMatch-RL adopts reinforcement learning in training.
@@
Recently, RAFT [107] estimates optical flow by iteratively updating a motion field with GRU. In MVS, several methods [40, 104, 103] also adopted this approach to enhance efficiency and flexibility. IterMVS [40] proposes a lightweight GRU-based probability estimator that encodes the per-pixel probability distribution of depth in its hidden state. CER-MVS [103] and Effi-MVS [104] both embed the RAFT module in a coarse-to-fine structure, which outputs a residual that is added to the previous depth. MaGNet [88] iteratively updates the Gaussian distribution of depth for each pixel with the matching scores. DELS-MVS [108] proposes Epipolar Residual Network to search for the corresponding point in the source image directly along the corresponding epipolar line and follows an iterative manner to narrow down the search space. Based on IterMVS [40], RIAV-MVS [109] iteratively refines index fields with a 3-level GRUs. IGEV-MVS [110] iteratively updates a disparity map regressed from geometry encoding cost volumes and pairs of correlation volumes.
3.5 Depth Estimation
3.5.1 Online MVS
@@ 3.5.2 Offline MVS
where $\mathbf{p}$ is the pixel coordinate, $d_{i}$ is $i$-th depth sample and $P(i,\mathbf{p})$ is the predicted depth probability. For coarse-to-fine methods [31, 38, 75], soft argmax is applied on each stage to regress the depth maps. On the other hand, some methods [35, 39] compute the expectation value of inverse depth samples since this sampling is more suitable for complex and large-scale scenes [35].
@@
Recently, Wang et al. [40] propose a hybrid strategy to combine regression and classification. Similarly, Peng et al. [111] first use classification to get the optimal hypothesis and then regress the proximity for it.
@@
3.6 Depth Refinement
Given that the raw depth estimation from MVS may be noisy, refinement is usually used to improve the accuracy. R-MVSNet [36] enforces multi-view photo-consistency to alleviate the stair effect and achieve sub-pixel precision. Point-MVSNet [97] uses PointFlow to refine the point cloud iteratively by estimating the residual between the depth of the current iteration and that of the ground truth. Fast-MVSNet [113] adopts an efficient Gauss-Newton layer to optimize the depth map by minimizing the feature residuals. PatchmatchNet [39] refines the final upsampled depth map with a depth residual network [114] and reference image feature.  [40, 104, 103] use the mask upsampling module from RAFT [107] to upsample and refine the depth map to full resolution by computing the weighted sum of depth values in a window based on the mask. Based on a coarse depth map, RayMVSNet [115, 116] aggregates multi-view image features with an epipolar transformer and use a 1D implicit field to estimate the SDF of the sampled points and the location of the zero-crossing point. GeoMVSNet [76] filters the depth map by geometry enhancement in the frequency domain. EPNet [117] uses a hierarchical edge-preserving residual learning module to refine multi-scale depth estimation with image context features.
3.7 Confidence Estimation
As discussed in Sec. 2.4, photometric confidence is important to filter out unreliable estimations during depth fusion. Following MVSNet [15], most offline MVS methods take the probability of the estimation [36, 37, 77] or the probability sum over several samples near the estimation [15, 31, 39] from the probability volume as confidence. In stereo matching, some methods learn confidence from disparity [118], RGB image [119, 120] or matching costs [121] and obtain confidence scores in $[0,1]$ interval. Motivated by this, some traditional MVS methods [122, 123] based on classical PatchMatch [106] propose to estimate the confidence with deep learning and use this to refine the results from PatchMatch. Recently, IterMVS [40] estimates confidence from the hidden state of a convolutional GRU. A 2D CNN followed by a sigmoid is applied to the hidden state to predict the confidence. DELS-MVS [108] feeds pixel-wise entropy of the partition probabilities, which is computed for each source image, into a confidence network to learn confidence. The confidence is used to guide the fusion of multiple depth maps.
3.8 Loss Function
3.8.1 Online MVS